pressure and volume of a gas gets doubled when a gas of definite mass is heated at 0°C. determine the final temperature of the gas

Grade:10

1 Answers

Eshan
5 years ago
From ideal gas equation,
PV=nRT
\implies \dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}
\implies T_2=4T_1=4\times 273K=1092K=819^{\circ}C